1. Convert each of the following to their respective Decimal, Octal, Hexadecimal and binary representation:

(a) (742)8

(b) (1011)2

(c) (47)10

(d) (3EAC)16

1
Expert's answer
2021-10-14T07:47:02-0400

(a) "(742)_8" is already in octal representation. So, we will change it to its Decimal, Hexadecimal and Binary form.

To Decimal

"(742)_8=7\times8^2+4\times 8^1+2\times 8^0\\
(742)_8=448+32+2=482_{10}."

To Hexadecimal form

"\def\arraystretch{1.5}
 \begin{array}{c|c}
 16&482 \\ \hline
 16&30~~r~~2\\
 16&1~~r~~E\\
&0~~r~~1
 
\end{array}"

"(742)_8=1E2_{16}"

To Binary form

"\def\arraystretch{1.5}
 \begin{array}{c|c}
 2&482 \\ \hline
 2&241~~r~~0\\
 2&120~~r~~1\\
2&60~~r~~0\\
2&30~~r~~0\\
2&15~~r~~0\\
 2&7~~r~~1\\
2&3~~r~~1\\
2&1~~r~~1\\
&0~~r~~1
\end{array}"

"(742)_8=111100010_{2}"


(b) "(1011)_2" is already in binary representation. So, we will change it to its Decimal, Hexadecimal and Octal form.

To Decimal

"(1011)_2=1\times2^3+0\times 2^2+1\times 2^1+1\times 2^0\\
(1011)_2=8+4+1=13_{10}."

To Hexadecimal form

"\def\arraystretch{1.5}
 \begin{array}{c|c}
 16&13 \\ \hline
 16&0~~r~~D\\
 
 
\end{array}"

"(1011)_2=D_{16}"

To Octal form

"\def\arraystretch{1.5}
 \begin{array}{c|c}
 8&13 \\ \hline
 8&1~~r~~5\\
 &0~~r~~1\\
\end{array}"

"(1011)_2=15_{8}"


(c) "(42)_{10}" is already in decimal representation. So, we will change it to its binary, Hexadecimal and Octal form.

To binary

"\def\arraystretch{1.5}
 \begin{array}{c|c}
 2&42 \\ \hline
 2&21~~r~~0\\
 2&10~~r~~1\\
2&5~~r~~0\\
2&2~~r~~1\\
2&1~~r~~0\\
 &0~~r~~1\\
\end{array}"

"(42)_{10}=101010_2"

To Hexadecimal form

"\def\arraystretch{1.5}
 \begin{array}{c|c}
 16&42 \\ \hline
 16&2~~r~~A\\
&0~~r~~2
 
 
\end{array}"

"(42)_{10}=2A_{16}"

To Octal form

"\def\arraystretch{1.5}
 \begin{array}{c|c}
 8&42\\ \hline
 8&5~~r~~2\\
 &0~~r~~5\\
\end{array}"

"(42)_{10}=52_{8}"


(d) "(3EAC)_{16}" is already in Hexadecimal representation. So, we will change it to its binary, decimal and Octal form.

To decimal

"(3EAC)_{16}=3\times16^3+E\times 16^2+A\times 16^1+C\times 16^0\\
(3EAC)_{16}=12288+3584+160+12=16044_{10}."


To Binary form

"(3EAC)_{16}=11111010101100_{2}"


To Octal form

"(3EAC)_{16}=37254_{8}"